OpenStudy (dedi):

If the discriminant of a quadratic equation is equal to -8, which statement describes the roots? A)There are two complex roots. B)There are two real roots. C)There is one real root. D)There is one complex root.

OpenStudy (campbell_st):

the discriminant is part of the general quadratic formula \[\Delta = b^2 - 4ac\] and the formula is \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] and the general quadratic formula is used to find the roots or zeros... so my question is this what is the value of \[\sqrt{-8}\]